if f(x)=2x-3and g(x)=x^2+1 fog(-1) (gof) (4)I thank you in advanced please help me here .and with this onefind (f-g) (x) f(x) =x^2 -2x-24 and g(x)=x^2 -x-30
Answer by stanbon(75887)

If f(x)=2x-3 and g(x)=x^2+1 fog(-1) (gof)(4)
---
fog(-1)
Step 1: find g(-1) = (-1)^2+1 = 1+1 = 2
So, f[g(-1)] = f(2)
Step 2: find f(2) = 2(2)-3 = (4-3) = 1
==================================================
find (f-g) (x) f(x) =x^2 -2x-24 and g(x)=x^2 -x-30
(f-g)x = f(x)-g(x)
Step 1: subtract g(x) from f(x)
= (x^2-2x-24) - (x^2-x-30)
Sept 2: simplify
= -2x+x -24+30
= -x + 6
==============
Cheers,
Stan H.
==============